wihtout graphing, decide whether the points would lie on the graph of the equation.

2x+4y=20
a. (6,2)
b. (2,4)
c. (10,4)

Answer by bucky(2189)

The points you are given are in the form (x, y) ... the first number in the parentheses
being the value of x and the number following the comma being the value in y. All you have
to do for each of the given points is to substitute the x value for x in the equation and
then also substitute the corresponding y value. If the equation still balances, then the
point is on the graph. So let's try this process for each of the three points.
First point is (6, 2) ... do x is 6 and y is 2. Substitute these two values for x and y in the
given equation as follows:
.
2x + 4y = 20
.
2(6)+ 4(2) = 20
.
do the two multiplications on the left side to get
.
12 + 8 = 20
.
And since the left side of the equation equals the right side, the point (6, 2) is on the graph.
.
Now the second point ... (2, 4). Substitute 2 for x and 4 for y:
.
2x + 4y = 20
.
2(2)+ 4(4) = 20
.
do the two multiplications on the left side to get
.
4 + 16 = 20
.
The left side of this equation equals the right side so the point (2, 4) is also on the graph.
.
Finally, the last point ... (10, 4) ....
.
2x + 4y = 20
.
2(10)+ 4(4) = 20
.
do the two multiplications on the left side to get
.
20 + 16 = 20
.
This time the left side does not equal the right side. Because of this you can say that the
point (10, 4) is NOT on the graph.
.
Hope this helps you to understand the problem and how to answer it.
.
.